- Service Observing
- At the telescope
- Remote Observing
- After the run
- AAO IT Services
- Gemini Office
- ASVO-AAT node
Occasionally, there may be reason to suspect that the performance of IRIS2 is abnormal, but there may be no clear way to prove this, or diagnose a problem. This Web page provides a set of "benchmarks", against which the current performance of the system can be compared. If you suspect a problem, please alert your support astronomer, the night assistant and/or the afternoon technician, and the instruement scientist.
The primary means of assessing the health of IRIS2's detector array is to execute the Array_Tests observing sequence, then allow ORAC-DR to process and analyse the results. The pipeline will alert the observer if either the read noise, or the dark current, is significantly outside of the expected ranges (DRM readout noise between 8 and 20 e-; dark current less than 1 e-/second). Note however that the recipe looks only at global statistics across the whole array; it might not recognise, for instance, a coherent noise pattern across one or more quadrants, which could indicate a more subtle problem. Thus, one should also take the time to examine each image individually from the Array_Tests sequence in the Skycat display. The image below is the typical result from a 6 x 10 sec DRM dark frame with the Mk2 science-grade array (post-May 2006) in the default orientation (N down, E to the right). Note in particular the group of hot pixels in the top right quadrant and its impact on the rows in which it lies (as well as the same rows in the quadrants below); low-level vertical artifacts in the lower-left quadrant due to hot pixels on the bottom edge; a large group of hot or dead pixels in the top-left corner; a smattering of persistent hot pixels across the array; and the elevated counts in the bottom 4 rows of each quadrant. These are all quite repeatable, and can be removed from your data with a matching dark frame.
In the event of poor array performance, the first remedy to try would be a Reset of the array controller. If only one or two quadrants appear to be suspect, then the afternoon technician may wish to swap the associated video cards.
Count rate suspiciously low
Assuming the array itself appears to be healthy, you may still suspect the throughput is not quite right. One possibility is that condensation may have formed on the front window of the IRIS2 fore-dewar, though this is now much less likely due to the double-insulation of the window, and the dome air-conditioning. A dead giveaway for moisture condensing on the front window is that it will appear as a dark patch in J- or H-band images of the sky, but will switch to being a bright patch in K-band due to thermal emission. If it does occur, it's time to get out the hair-drier and hand mirror.
If you're still not sure if things look right, compare a raw on-sky exposure with >10K ADU per pixel with this Ks-band sky image (N down, E to right). The flatfield structure seen in J- or H-band is virtually identical to this.
If the count rates are still suspiciously low, and you're certain it really is photometric out there, then go ahead and observe an infared standard, and compare your derived zero-point to the tabulated IRIS2 sensivities, taking account of the net exposure time given by the EXPOSED keyword. There are several ways you might do this: using the imexam task or apphot package within IRAF, or the Image-Analysis -> Aperture Photometry -> Results in magnitudesGAIA utility, for example.
Focus and Image Quality
Don't forget that there are two mechanisms for controlling the focus: the detector is focused onto the focal plane by moving the array translator, and then the telescope is focused onto the focal plane by moving the secondary mirror. There is a known change in IRIS2 focus with wavelength/filter, but this is almost negligible compared with the amount of array translation required to make any significant difference to the delivered image quality. If the images appear out of focus, but still circular, then the first thing to suspect is the array translator focus. The Observing Sequence just executed may contain an inappropriate detector focus command, or the default Focus value of 0 may have been set inadvertently while commanding a manual configuration from the spectrograph GUI.
There is also a known problem with astigmatic images produced by the f/8 secondary mirror at some extreme zenith distances, particularly when working in the far south and when the seeing is sub-arcsecond. The signature of this is elliptical images, the major axes of which flip through 90o when the secondary mirror is moved either side of the nominal focus position. The cause is thought to be slop in the mounting of the oversize f/8 secondary mirror. The only known cure currently is to go and observe in another area of the sky, or wait for the seeing to degrade to such an extent that the astigmatism no longer dominates!
Wheels don't seem to move
If you have reason to suspect that one or more wheels inside the dewars are stuck, or a wheel won't home properly, have someone go out into the Cass cage and hold their ear up against the dewar wall when a wheel movement is commanded. The sound of the motors driving, and the wheel moving should be fairly apparent, as will the "clunk" when the drive worm is engaged or disengaged. Checking photometric zero-points, by observing standard stars as described above, may be the best way to establish which filter is actually in the beam.